OpenStudy (anonymous):

If a=\(1-\frac{ 1 }{ 1-b }\), then b=?

hartnn (hartnn):

in terms of a ?

OpenStudy (anonymous):

YES

hartnn (hartnn):

so, how did u start ? u got anything ?

hartnn (hartnn):

1/(1-b)=1-a could you get this far ?

OpenStudy (anonymous):

b= (-a)/(-a-1) i don't think this is the answer lolz

hartnn (hartnn):

ok, 1/(1-b)=1-a got this much ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

good, now you can flip the fraction, can you ? or should i give an example ?

OpenStudy (anonymous):

b=1-(1/(1-a))?

hartnn (hartnn):

correct! good, now simplify :)

OpenStudy (anonymous):

the multiple choice question has that answer, thank you very much

hartnn (hartnn):

oh, nice :) welcome ^_^